OpenStudy (anonymous):

use the separation of variables to solve the initial value problem: dy/dx = x/y when y=2 and x = 1

OpenStudy (anonymous):

no that is exactly what was given to me

OpenStudy (displayerror):

That is a separable differential equation, so manipulate the equation such that all of the 'y' terms are on one side and all of the 'x' terms are on the other side: Your equation: \[\frac{dy}{dx} = \frac{x}{y}\] \[\int{y \ dy} = \int{x \ dx}\] It's an indefinite integral, so remember that there is a constant \(C\). You have to use your initial condition to solve for that constant.

OpenStudy (solomonzelman):

oh, yeah, I see what they are asking you. \(\large\color{black}{ \displaystyle \frac{dy}{dx} =\frac{x}{y} }\) \(\large\color{black}{ \displaystyle y\frac{dy}{dx} =x }\) \(\large\color{black}{ \displaystyle \color{blue}{\int\limits_{~}^{~}}y\frac{dy}{dx} \color{blue}{dx}=\color{blue}{\int\limits_{~}^{~}}x\color{blue}{dx} }\)

OpenStudy (solomonzelman):

this is how it's done technically.

OpenStudy (solomonzelman):

I integrated with respect to x on both sides, not dx's on the left side cancel

OpenStudy (solomonzelman):

then integrate each side, and to find the +C plug in your values of y and x.

OpenStudy (solomonzelman):

you can combine all your +C into one +C on the right side.

OpenStudy (anonymous):

Okay I'm still a little sketchy on taking integral and the notation for that. I know it involves antiderivative but can we go through finding the integral?

OpenStudy (solomonzelman):

Ok, what is \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}x~dx}\) ?

OpenStudy (solomonzelman):

(just apply the power rule)

OpenStudy (anonymous):

\[1/2 x^{2}\]

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle\int\limits_{~}^{~}x~dx=\frac{1}{2}x^2+C}\)

OpenStudy (solomonzelman):

And the left side, (after you cancel the dx's on both sides) is: \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}y~dx}\) And, \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}y~dx=\frac{1}{2}y^2+C}\) (integrating y with respect to y, and integrating x with respect to x -- is a same thing)

OpenStudy (anonymous):

Okay add the constant. That makes sense. The rules for y are still confusing

OpenStudy (solomonzelman):

for y it is just the same thing

OpenStudy (anonymous):

okay so there is nothing special about the y at all? Alright then. So now you fill in x and y that were given to me and solve for c

OpenStudy (solomonzelman):

yes, in this case there is nothing special about y.

OpenStudy (solomonzelman):

(unlike, as you know adding a chain dy/dx for a derivative of y -- nothing like this when integrating)

OpenStudy (anonymous):

Right that was what was confusing me on that part

OpenStudy (solomonzelman):

and, then \(\large\color{slate}{\displaystyle\frac{1}{2}y^2+C_1=\frac{1}{2}x^2+C_2}\) \(\large\color{slate}{\displaystyle\frac{1}{2}y^2=\frac{1}{2}x^2+C_2-C_1}\) \(\large\color{slate}{\displaystyle\frac{1}{2}y^2=\frac{1}{2}x^2+C_3}\) C(3) -- still some arbitrary constant.

OpenStudy (solomonzelman):

So your function just is: \(\large\color{slate}{\displaystyle\frac{1}{2}y^2=\frac{1}{2}x^2+C}\)

OpenStudy (solomonzelman):

you can divide by 1/2 everywhere

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle y^2=x^2+C}\) multiplying an arbitrary constant by 2 still leaves it an arbitrary constant

OpenStudy (solomonzelman):

then: \(\large\color{slate}{\displaystyle 2^2=1^2+C}\) >> \(\large\color{slate}{\displaystyle C=3}\)

OpenStudy (solomonzelman):

hope this is making sense

OpenStudy (anonymous):

Okay then my last question. Does it matter what form the function is in before plugging in? As in do I need to get the y by itself or anything or can I just start plugging in the variables I was given?

OpenStudy (solomonzelman):

actually, I shouldn't multiply by 2 before finding C.

OpenStudy (solomonzelman):

well, or I can by I should say: \(\large\color{slate}{\displaystyle y^2=x^2+2C}\)

OpenStudy (solomonzelman):

because this C that we got up to, is the actual vertical shift of the function.

OpenStudy (solomonzelman):

so by multiplying times 2 everything instead of C I doubled the value of C (and I shouldn't do that)

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle y^2=x^2+2C}\) \(\large\color{slate}{\displaystyle 2^2=1^2+2C}\) \(\large\color{slate}{\displaystyle 3=2C}\) C=3/2

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle y^2=x^2+3}\)

OpenStudy (anonymous):

Alright I had to leave for a while. But I got it thank you SO MUCH!!!

OpenStudy (solomonzelman):

surely, yw!